My prealgebra class is studying integers and evaluating expressions- I broke my leg last week and was not in class so I'm kinda of lost on this, can you help? here's my problem 1-3-(-2)=1+(a)+2
=-2+(b)
=C

Answer by Alan3354(69443)

1-3-(-2)=1+(a)+2
=-2+(b)
=C
----------
Is it one problem? Or 3?
1-3-(-2)=1+(a)+2
-2 -(-2)=1+(a)+2
-2 + 2 = 3 + a
a + 3 = 0
a = -3